# WikiAutoGen: Towards Multi-Modal Wikipedia-Style Article Generation

### v1

@@ Abstract
Knowledge discovery and collection are intelligence-intensive tasks that traditionally require significant human effort to ensure high-quality outputs. Recent research has explored multi-agent frameworks for automating Wikipedia-style article generation by retrieving and synthesizing information from the internet. However, these methods primarily focus on text-only generation, overlooking the importance of multimodal content in enhancing informativeness and engagement. In this work, we introduce WikiAutoGen, a novel system for automated multimodal Wikipedia-style article generation. Unlike prior approaches, WikiAutoGen retrieves and integrates relevant images alongside text, enriching both the depth and visual appeal of generated content. To further improve factual accuracy and comprehensiveness, we propose a multi-perspective self-reflection mechanism, which critically assesses retrieved content from diverse viewpoints to enhance reliability, breadth, and coherence, etc. Additionally, we introduce WikiSeek, a benchmark comprising Wikipedia articles with topics paired with both textual and image-based representations, designed to evaluate multimodal knowledge generation on more challenging topics. Experimental results show that WikiAutoGen outperforms previous methods by 8%-29% on our WikiSeek benchmark, producing more accurate, coherent, and visually enriched Wikipedia-style articles. We show some of our generated examples in https://wikiautogen.github.io/
@@ 1 Introduction
Knowledge discovery and content generation are essential for organizing and disseminating information, but they remain time-consuming and intelligence-intensive, requiring significant human effort to collect, structure, and verify information. With the advent of large-scale AI models like large language model (LLM) [27, 10, 43, 2, 5], there is growing potential to automate knowledge collection, synthesis, and organization in a more efficient and scalable manner [17, 33]. Such automation not only accelerates knowledge discovery but also enhances accessibility, making information more readily available and up to date.
Recently, several methods, such as Storm [33] and Co-Storm[17], have been proposed to automate Wikipedia-style article generation. While these approaches can produce Wikipedia-like content, they still face key limitations: (1) they are restricted to text-only generation, failing to incorporate multimodal content like relevant images; (2) the generated articles often lack breadth, depth, and reliability, reducing their overall informativeness and credibility.
To address these challenges, we introduce WikiAutoGen, a multi-agent framework designed to automatically generate high-quality, multimodal Wikipedia-style articles. Unlike prior works, WikiAutoGen can directly search both textual and visual content and generate multi-modal content (see Fig 1), enhancing the informativeness and engagement of generated articles. Additionally, we propose a novel multi-perspective self-reflection module, which enables the system to self-regulate, refine, and critically evaluate its generated content. This mechanism enhances the reliability, depth, and breadth of the information by encouraging iterative improvement and multi-source validation.
To advance the development and evaluation of multimodal knowledge generation, we introduce WikiSeek, a new benchmark designed to tackle challenging topics represented through both images and text. Existing benchmarks primarily focus on text generation or cover only straightforward topics (see Table. 1). In contrast, WikiSeek is multimodal and specifically targets more complex subjects with limited coverage on Wikipedia, making it significantly harder for current methods to retrieve and synthesize comprehensive information. This increases the challenge of content generation, pushing models to explore deeper, enhance their retrieval capabilities, and improve their ability to handle underexplored subjects.
@@
Table 1: Comparison of WikiSeek with existing benchmarks. Modalities are indicated by [Uncaptioned image] (text) and [Uncaptioned image] (images). Difficulty levels are categorized based on the average number of characters in corresponding Wikipedia pages: difficult ($<$500 characters), medium (500–2000 characters), and easy ($>$2000 characters).
Extensive experiments demonstrate that WikiAutoGen significantly outperforms existing methods in generating high-quality textual and visual content. We evaluated text quality across nine key dimensions and image quality across four essential criteria. Experimental results show that WikiAutoGen achieves improvements ranging from 8% to 29% in textual evaluations and from 11% to 14% in image evaluations compared to baseline methods for the input topics of text-only, image-only and image-text.
@@
  • We propose WikiAutoGen, a multimodal article generation framework capable of synthesizing comprehensive content by effectively integrating textual and visual inputs. Evaluations on the WikiSeek benchmark demonstrate that WikiAutoGen achieves significant performance improvements of 8%–29% over previous best model on textual generation.
@@
  • We develop a novel multi-perspective self-reflection module, enhancing the readability, informativeness, and coherence of generated articles by providing iterative feedback from diverse viewpoints, such as reader, writer, and editor.
2 Realted Work
@@
Automatic Expository Writing. LLMs have demonstrated significant practical value in automatic expository writing, generating human-like articles, such as Wikipedia-style surveys [38, 20, 49, 26, 4, 45]. While LLMs excel in traditional Natural Language Generation (NLG) tasks, producing long-form, well-structured, coherent, and logically organized text remains a persistent challenge [31, 12, 39, 34, 44]. To address this, [37] introduced domain-specific content keywords and progressively refined them into complete passages through multiple stages. Shao et al. [33, 17] highlighted the crucial role of pre-writing strategies, identify them as key a factor in writing systems that significantly enhance the quality of generated articles. Recently, some automatic writing systems expanded their knowledge boundaries through mindmaps and tree-based methods [3, 23, 34, 48]. While these iterative content planning methods effectively leverage widely accessible information for common topics, they prove largely ineffective in less-explored or niche domains, which have been overlooked [13, 50]. Meanwhile, existing methods are limited to generating text-based articles and fail to incorporate other modalities, such as visual elements. This inherent constraint in data input inevitably leads to information loss and incompleteness, reducing the article readability and making knowledge acquisition more challenging. Conversely, this work integrates visual content and retrieves knowledge across multiple modalities, allowing visual information to complement textual content by capturing details that may otherwise be overlooked.
@@
Self-reflection. Recent progress in optimizing LLMs through self-reflection mechanisms is significant. The core is to let models analyze and refine outputs with auto-generated feedback. Existing approaches construct feedback sources from three angles: (1) The LLM conducts iterative self-evaluation [35]; (2) A separately trained critic module provides specialized feedback [14]; (3) External knowledge from sources like Wikipedia and browsers is integrated [25, 1]. Notably, REFINER [29] proved that a trained critic module can enhance reasoning without fine-tuning the reasoning module, supporting feedback mechanism optimization. Meanwhile, some methods [9, 55] created a feedback mechanism using error-type templates and context-hypothesis mappings. Recent studies [30, 47, 36, 46, 53] focus on LLMs’ in-context learning. They design prompt templates to help models generate good feedback from historical outputs or patterns. Notably, the multi-perspective self-reflection method by [51], which relies on a Navigator-Reasoner heuristic interaction, is limited to LLM-based discussion without new knowledge acquisition, leading to low information richness and verifiability. In contrast, our approach enhances multi-perspective self-reflection with multi-web search, addressing complex topic exploration and retrieval challenges while enabling multi-dimensional control over topic and article quality.
Retrieval-Augmented Generation. Retrieval-Augmented Generation(RAG) has shown great potential for knowledge-intensive and information-seeking tasks, which has been widely used to improve the generation quality of language models by incorporating external knowledge [8, 52, 7, 1, 15, 18]. Although RAG has been extensively explored in language domains and its application in multi-modal contexts is advancing, despite the great potential of text-based RAG, it can only handle textual inputs and cannot utilize the rich information in multi-modal data such as images and videos as noted in [41, 42, 40, 56, 54]. However, the applications of multi-modal RAG (MRAG) remain largely task-specific [24, 22]. Also, there are multimodal RAG approaches, and no comprehensive framework exists for integrating MRAG into automatic expository writing. Moreover, both single- and multi-stage approaches struggle with multi-modal ambiguity, such as uncertain text-image associations, leading to suboptimal responses.
@@ 3.1 Task Definition
The task is to automatically generate a multi-modal Wikipedia-style article. Given a topic represented as text $T$, an image $I$, or a combination of both ($T,I$), the goal is to generate an article $A$ that cites a list of references $R$, incorporating relevant and supporting knowledge. This capability is particularly crucial in domains such as investigative journalism, scientific research, and market analysis, where reliable and well-sourced information is essential.
@@ 3.2 WikiSeek Construction
A key challenge in this task is the lack of an appropriate benchmark that includes multimodal topics. Existing benchmarks primarily focus on textual topics, failing to capture the complexities of multimodal content generation. To bridge this gap, we introduce WikiSeek, a new benchmark designed to evaluate more challenging topics that incorporate both text and images. WikiSeek provides a rigorous evaluation framework, enabling a more comprehensive assessment of multimodal knowledge generation. In the following sections, we outline the construction process of this benchmark in detail.
Figure 2: Overview of WikiAutoGen, our multimodal framework for Wikipedia-style article generation. The pipeline includes three main stages: (1) an Outline Proposal module that structures the article outline based on the multimodal topic input (image and text); (2) a Textual Article Writing module involving persona generation, multi-agent collaborative exploration, and article drafting; and (3) a Multimodal Article Writing module that incorporates relevant images through positioning proposals, retrieval, selection, and final polishing. The entire generation process is enhanced by a Multi-Perspective Self-Reflection module, leveraging supervisory and agent-specific feedback (writer, reader, editor) to iteratively improve article quality in terms of coherence, readability and engagement, etc
Benchmark construction pipeline. Our WikiSeek benchmark is designed with two key objectives: 1) To evaluate multi-modal article generation, where both the querying topics and output consist of text and images. 2) To focus on more challenging topics that are underexplored on Wikipedia, where limited relevant information is available, making retrieval and synthesis more difficult.
@@
However, some topics, such as “1997 in Japan” and “.bh”, are overly broad or too abstract, making them unsuitable for evaluation. To address this, we manually verify all selected topics and remove those that lack meaningful content or pose evaluation challenges.
After this filtering process, we obtain a final set of 300 topics, evenly distributed across three difficulty levels (hard, very hard, and extremely hard), with 100 topics per level. These topics are represented in one of three formats: text-only, image-only, or a combination of text and images.
@@ 4 Method
Writing a high-quality multimodal Wikipedia-style article usually requires a multi-agent system that effectively breaks down the process into distinct stages: e.g., proposing an outline, retrieving relevant materials from the internet, and synthesizing the final article. Beyond these stages, maintaining quality necessitates collaboration across multiple roles, ensuring the article is well-structured, accurate, and engaging. To address these challenges, we introduce WikiAutoGen framework. This framework facilitates structured collaboration among specialized agents, enabling comprehensive topic exploration, multi-modal content generation, and coherent content generation. In the following sections, we provide a detailed breakdown of WikiAutoGen and its core components.
@@ 4.1 WikiAutoGen Framework
Our WikiAutoGen framework consists of several key components that work collaboratively to generate high-quality multimodal Wikipedia-style articles, as illustrated in Fig. 2. Each module plays a distinct role in the article generation process: 1). Outline Proposal Module. This module converts the given text and image topics into structured outline proposals, laying the foundation for content organization. 2). Textual Article Writing Module. This stage involves multiple sub-components, including a persona generator, a multi-agent discussion system, and the article generation process, ensuring the content is well-structured and contextually rich. 3). Multi-Perspective Self-Reflection Module. This component evaluates the generated text from multiple viewpoints, including those of a writer, reader, and editor, providing constructive feedback to refine and enhance the article. 4). Multimodal Article Writing Module. This final stage integrates visual content, consisting of image positioning proposals, image retrieval, image selection, and multimodal refinement, ensuring a cohesive and well-balanced article presentation. These components collectively enable WikiAutoGen to produce high-quality and multi-modal Wikipedia-style articles.
@@ 4.2 Outline Proposal
Given a multimodal topic, the first step is to interpret the input and generate topic-related outlines to guide knowledge exploration and information retrieval. We achieve this by leveraging a large language model (LLM) [16] and external search tools.
For text-based topics, the LLM analyzes the input, identifies relevant subtopics, and generates a structured outline to facilitate knowledge exploration. For image-based topics, we utilize Google Vision Search to retrieve metadata, including descriptions and contextual information. Named entity recognition (NER) [11] is then applied to extract the top 10 most frequent entities as query keywords. These keywords, along with the original topic, are then fed into the LLM to generate a structured outline.
In the case of image-text topics, we combine insights from both modalities by extracting subtopics from the text and retrieving image metadata. The LLM then refines the topic and synthesizes a cohesive outline, integrating textual and visual information to guide comprehensive knowledge retrieval.
@@ 4.3 Textual Article Writing
Persona generator. Given a draft outline, the LLM generates n distinct roles relevant to the topic (where n is a customizable parameter, $n\geq 1$), each acting as an independent agent. The LLM assigns specific objectives to each agent based on their role. These agents are equipped with the ability to connect to a search engine, allowing them to retrieve relevant information and contribute to a more comprehensive and well-supported article.
Multi-agent knowledge exploration. The knowledge exploration stage involves a fixed agent, the asker, and n LLM-generated agents assigned specific roles. The asker iterates through the outline, posing targeted questions, while the other agents search the internet for relevant information. They then share their findings, discuss them, and refine their understanding. During the discussion, they also interact with the multi-perspective self-reflection module, which provides feedback and improvement advise from a writer’s perspective on reliability, engagement, consistency and informativeness. This iterative process ensures a well-rounded knowledge base before proceeding to article generation.
Article generation. After gathering web knowledge, the next step is to summarize the collected content and generate the textual article using an LLM-based writing agent. Once the initial draft is produced, the agent iteratively send each generated section to the multi-perspective self-reflection module for feedback and then refine each paragraph. This self-reflection module evaluates the article from a writer’s perspective, providing suggestions to enhance coherence and readability. The writing agent then incorporates these refinements, iteratively improving the text to produce the final article.
@@
Writing a high-quality Wikipedia-style article requires addressing multiple aspects, including topic consistency, readability, engagement, and informativeness, to provide an optimal reading experience. Therefore, we introduce a multi-perspective self-reflection module that systematically evaluates and refines content across these dimensions. This module take four distinct viewpoints and assess the article from seven perspectives.
Perspectives. Our multi-perspective self-reflection focus on the seven key criteria to improve the paper writing quality. They include reliability, engagement, informativeness, coherence, readability, consistency and helpfulness. We provide detail explanation for them in the Appendix.
Supervisor viewpoint. The supervisor assesses whether the generated content fully addresses the questions posed by the asker, evaluates the article’s depth, breadth, and coherence, and reviews the effectiveness of the multi-agent discussion. Additionally, it evalutes whether the generated content aligns with the topic and proposed outlines. Based on these criteria, the supervisor provides an evaluation and passes the feedback to the next role. Depending on the specific needs, the next role can be the writer, reader, or editor.
@@
Table 2: Comparison of article generation performance for textual content. We evaluate on textual generation for content quality, informativeness, reliability, and engagement, under three input modalities (Text-only, Image-only, and Image-Text) on our WikiSeek benchmark.
Writer viewpoint. From the writer’s viewpoint, the primary focus is on the multi-agent knowledge exploration and article generation stages. The writer evaluates whether the generated content maintains coherence, ensures engagement, verifies factual accuracy, and upholds logical consistency. Based on these assessments, the writer provides targeted improvement suggestions. For instance, to enhance coherence, it may recommend rearranging sentences or adding transitional words and phrases. To improve readability, it might suggest simplifying complex concepts. Finally, the writer responds with a set of refined suggestions.
@@
Editor viewpoint. After inserting images into the generated article, there may still be discrepancies or inconsistencies between the visual content and the corresponding textual descriptions. To address this, our framework sends the images along with their related text segments to the multi-perspective self-reflection module. This module evaluates the alignment and coherence between the images and their accompanying texts from an editorial viewpoint. It provides targeted suggestions, such as refining the image captions, adjusting image placement, or enhancing textual explanations to better reflect visual content. This step ensures improved relevance, coherence, and readability.

### v2

@@ Abstract
Knowledge discovery and collection are intelligence-intensive tasks that traditionally require significant human effort to ensure high-quality outputs. Recent research has explored multi-agent frameworks for automating Wikipedia-style article generation by retrieving and synthesizing information from the internet. However, these methods primarily focus on text-only generation, overlooking the importance of multimodal content in enhancing informativeness and engagement. In this work, we introduce WikiAutoGen, a novel system for automated multimodal Wikipedia-style article generation. Unlike prior approaches, WikiAutoGen retrieves and integrates relevant images alongside text, enriching both the depth and visual appeal of generated content. To further improve factual accuracy and comprehensiveness, we propose a multi-perspective self-reflection mechanism, which critically assesses retrieved content from diverse viewpoints to enhance reliability, breadth, and coherence, etc. Additionally, we introduce WikiSeek, a benchmark comprising Wikipedia articles with topics paired with both textual and image-based representations, designed to evaluate multimodal knowledge generation on more challenging topics. Experimental results show that WikiAutoGen outperforms previous methods by 8%-29% on our WikiSeek benchmark, producing more accurate, coherent, and visually enriched Wikipedia-style articles. Our code and examples are available at https://wikiautogen.github.io/
@@ 1 Introduction
Knowledge discovery and content generation are essential for organizing and disseminating information, but they remain time-consuming and intelligence-intensive, requiring substantial human effort to collect, structure, and verify information. With the advent of large-scale AI models like large language models (LLMs) [22, 8, 35, 2, 5], there is growing potential to automate knowledge collection, synthesis, and organization in a more efficient and scalable manner [14, 28]. Such automation not only accelerates knowledge discovery but also enhances accessibility, making information more readily available and up to date.
Recently, several methods, such as Storm [28] and Co-Storm [14], have been proposed to automate Wikipedia-style article generation. While they can produce Wikipedia-like content, they still face key limitations: (1) they are limited to text-only generation and lack the ability to incorporate multimodal content such as relevant images; (2) the generated articles often lack breadth, depth, and reliability, reducing their overall informativeness and credibility.
To address these challenges, we introduce WikiAutoGen, a multi-agent framework designed to generate high-quality, multimodal Wikipedia-style articles automatically. Unlike prior works, WikiAutoGen can directly search both textual and visual information and generate multimodal content (see Figure 1), enriching article content with relevant and diverse modalities. Additionally, we propose a novel multi-perspective self-reflection module, which enables the system to self-regulate, refine, and critically evaluate its generated content. This mechanism enhances the reliability, depth, and breadth of the information by encouraging iterative improvement and multi-source validation.
To advance the development and evaluation of multimodal knowledge generation, we introduce WikiSeek, a new benchmark designed to tackle challenging topics comprising both textual and visual components. Existing benchmarks primarily focus on text generation or cover only straightforward topics (see Table 1). In contrast, WikiSeek is multimodal and specifically targets more complex subjects with limited coverage on Wikipedia, making it significantly harder for current methods to retrieve and synthesize comprehensive information. This increases the challenge of content generation, pushing models to explore deeper, enhance their retrieval capabilities, and improve their ability to handle underexplored subjects.
@@
Table 1: Comparison of WikiSeek with existing benchmarks. Modalities are indicated by [Uncaptioned image] (text) and [Uncaptioned image] (images). Difficulty levels are categorized based on the average number of characters in corresponding Wikipedia pages: difficult (<<500 characters), medium (500–2000 characters), and easy (>>2000 characters).
Extensive experiments demonstrate that WikiAutoGen significantly outperforms existing methods in generating high-quality textual and visual content. We evaluated text quality across nine key dimensions and image quality across four essential criteria. Experimental results show that WikiAutoGen outperforms prior methods by 8%–29% in textual quality and 11%–14% in image quality, demonstrating consistent gains across the input topics of text-only, image-only, and image-text tasks.
@@
  • We propose WikiAutoGen, a multimodal article generation framework that synthesizes comprehensive content by effectively integrating textual and visual inputs. On WikiSeek, WikiAutoGen achieves 8%–29% improvements over the previous best model in textual generation.
@@
  • We develop a novel multi-perspective self-reflection module, which iteratively enhances readability, informativeness, and coherence by incorporating feedback from diverse roles, including reader, writer, and editor.
2 Related Work
@@
Automatic Expository Writing.  LLMs have shown strong performance in automatic expository writing, particularly in generating Wikipedia-style articles [33, 17, 41, 21, 4, 37]. Despite their strength in traditional Natural Language Generation (NLG), LLMs still struggle to produce long-form text that is coherent and logically structured [26, 10, 34, 29, 36]. To address this, [32] proposed using domain-specific keywords that are progressively refined into full passages through multi-stage generation. Notably, Shao et al. [28, 14] highlighted the crucial role of pre-writing strategies, identifying them as a key factor in improving article quality. Recently, some automatic writing systems expanded their knowledge boundaries through mindmaps and tree-based methods [3, 19, 29, 40]. However, while these iterative content planning methods effectively leverage widely accessible information for common topics, they remain largely ineffective in less-explored or niche domains due to the scarcity of structured prior knowledge [11, 42]. Meanwhile, existing methods are limited to generating text-based articles and fail to incorporate other modalities, such as visual elements. This inherent constraint in data input inevitably leads to incomplete information and reduced readability, making knowledge acquisition more challenging. Conversely, our WikiAutoGen is the first multimodal writing system that integrates visual content and retrieves knowledge across multiple modalities, allowing visual information to complement textual content by capturing details that may otherwise be overlooked.
@@
Self-reflection. Recent progress in optimizing LLMs through self-reflection mechanisms is significant. The core idea is to enable models to analyze and refine their outputs through self-generated feedback. Existing approaches construct feedback sources from three strategies: (1) The LLM conducts iterative self-evaluation [30]; (2) A separately trained critic module provides specialized feedback [12]; (3) External knowledge from sources like Wikipedia and browsers is integrated [20, 1]. Specifically, REFINER [24] demonstrated that a trained critic module can enhance reasoning without fine-tuning the reasoning module, supporting feedback mechanism optimization. Further, some methods [7, 45] introduced feedback mechanisms based on error-type templates and context-hypothesis mappings. Recent studies [25, 39, 31, 38, 44] focus on LLMs' in-context learning. They design prompt templates to help models generate feedback from historical outputs or patterns. However, the multi-perspective self-reflection method proposed by [43], which relies on a Navigator-Reasoner heuristic interaction, is limited to closed-loop LLM discussions without external knowledge acquisition, resulting in restricted information richness and verifiability. In contrast, our approach enhances multi-perspective self-reflection with multi-web search, addressing complex topic exploration and retrieval challenges while enabling multi-dimensional control over topic and article quality.
@@ 3.1 Task Definition
Given a topic as text ($T$), image ($I$), or a combination of both ($T,I$), the objective is to generate a Wikipedia-style article ($A$) that integrates relevant knowledge and is supported by verifiable references ($R$). This task is particularly important in domains such as investigative journalism, scientific research, and market analysis, where the generation of accurate, well-sourced content is essential.
@@ 3.2 WikiSeek Construction
A key challenge in this task is the lack of a suitable benchmark that effectively encompasses multimodal topics. Existing benchmarks, however, remain largely text-centric, thus failing to adequately reflect the complexity of multimodal content generation. To address this limitation, we introduce WikiSeek, a new benchmark specifically designed to evaluate more challenging topics that incorporate both text and images. WikiSeek establishes a robust evaluation framework, enabling a more comprehensive and reliable assessment of multimodal knowledge generation in practice. In the following sections, we detail the construction process of this benchmark.
Figure 2: Overview of WikiAutoGen, our multimodal framework for Wikipedia-style article generation. The pipeline includes three main stages: (1) an Outline Proposal module that structures the article outline based on the multimodal topic input (image and text); (2) a Textual Article Writing module involving persona generation, multi-agent collaborative exploration, and article drafting; and (3) a Multimodal Article Writing module that incorporates relevant images through positioning proposals, retrieval, selection, and final polishing. The entire generation process is enhanced by a Multi-Perspective Self-Reflection module, leveraging supervisory and agent-specific feedback (writer, reader, editor) to iteratively improve article quality in terms of coherence, readability, and engagement, etc.
Benchmark construction pipeline. Our WikiSeek benchmark is designed with two key objectives: (1) to evaluate multimodal article generation where both the input and output include text and images; and (2) to target underexplored topics on Wikipedia that present greater challenges for retrieval and synthesis.
@@
However, some topics, such as ``1997 in Japan" and ``.bh", are either overly general or semantically underspecified, making them unsuitable for evaluation. To address this, we manually verify all selected topics and remove those that lack meaningful content or pose evaluation challenges. After this filtering process, we obtain a final set of 300 topics, evenly distributed across three difficulty levels (hard, very hard, and extremely hard), with 100 topics per level. These topics are represented in one of three formats: text-only, image-only, or a combination of both.
@@ 4 Method
Writing high-quality multimodal Wikipedia-style articles usually requires a coordinated multi-agent system that effectively breaks down the process into distinct stages, including outline generation, web-based material retrieval, and article synthesis. Beyond these stages, maintaining quality necessitates collaboration across roles, ensuring the article is well-structured, accurate, and engaging. To address these challenges, we propose WikiAutoGen framework. This framework facilitates structured collaboration among specialized agents, enabling comprehensive topic exploration, multi-modal content generation, and coherent generation. In the following sections, we provide a detailed breakdown of WikiAutoGen and its core components.
@@ 4.1 WikiAutoGen Framework
Our WikiAutoGen framework consists of several key components that work collaboratively to generate high-quality multimodal Wikipedia-style articles, as illustrated in Figure 2. Each module plays a distinct role in the article generation process: 1). Outline Proposal Module. This module converts the given text and image topics into structured outline proposals, laying the foundation for content organization. 2). Textual Article Writing Module. This stage involves multiple sub-components, including a persona generator, a multi-agent discussion system, and the article generation process, ensuring the content is well-structured and contextually rich. 3). Multi-Perspective Self-reflection Module. This component evaluates the generated text from multiple viewpoints, including those of a writer, reader, and editor, providing constructive feedback to refine and enhance the article. 4). Multimodal Article Writing Module. This final stage integrates visual content, consisting of image positioning proposals, image retrieval, image selection, and multimodal refinement, ensuring a cohesive and well-balanced article presentation. These components collectively enable WikiAutoGen to produce high-quality and multi-modal Wikipedia-style articles.
@@ 4.2 Outline Proposal
Given a multimodal topic, the first step is to interpret the input and generate topic-related outlines to guide knowledge exploration and information retrieval. We achieve it by leveraging LLMs [13] and external search tools.
For text-based topics, the LLM analyzes the input, identifies relevant subtopics, and generates a structured outline to facilitate knowledge exploration. For image-based topics, we utilize Google Vision Search to retrieve metadata, including descriptions and contextual information. Named entity recognition (NER) [9] is then applied to extract the top 10 most frequent entities as query keywords. These keywords, combined with the original topic, are then fed into the LLM to generate a well-structured outline. In the case of image-text topics, we combine insights from both modalities by extracting subtopics from the text and retrieving image metadata. The LLM then refines the topic and synthesizes a cohesive outline, integrating textual and visual information to guide comprehensive knowledge retrieval.
@@ 4.3 Textual Article Writing
Persona generator. Given a draft outline, the LLM generates $n$ distinct personas relevant to the topic (where n is a customizable parameter, $n\geq 1$), each acting as an independent agent. The LLM assigns specific objectives to each agent based on their role. These agents are equipped with access to external web search tools and contribute to a more comprehensive and well-supported article.
Multi-agent knowledge exploration. The knowledge exploration stage involves a fixed agent, the asker, and n LLM-generated agents who are assigned specific roles. The asker iterates through the outline, posing targeted questions, while the other agents search the internet for relevant information. They then share findings, discuss them, and refine their understanding. During the discussion, they also interact with the multi-perspective self-reflection module, which provides feedback and improvement advice from a writer's perspective on reliability, engagement, consistency, and informativeness. This iterative process ensures a well-rounded knowledge base before article generation.
Article generation. After gathering web knowledge, the next step is to summarize the collected content and generate the textual article using an LLM-based writing agent. Once the initial draft is produced, the agent iteratively sends each generated section to the multi-perspective self-reflection module for feedback and refines each paragraph. This module evaluates the article from a writer’s perspective, providing suggestions to enhance coherence and readability. The writing agent incorporates these refinements, progressively improving the text to produce the final article.
@@
Writing a high-quality Wikipedia-style article requires addressing multiple aspects, including topic consistency, readability, engagement, and informativeness, to provide an optimal reading experience. Therefore, we introduce a multi-perspective self-reflection module that systematically evaluates and refines content across these dimensions. This module takes four distinct viewpoints and assesses the article from seven perspectives.
Perspectives. Our multi-perspective self-reflection focuses on the seven key criteria to improve the paper writing quality. They include reliability, engagement, informativeness, coherence, readability, consistency and helpfulness. We provide a detailed explanation for them in the Appendix.
Supervisor viewpoint. The supervisor assesses whether the generated content fully addresses the questions posed by the asker, evaluates the article’s depth, breadth, and coherence, and reviews the effectiveness of the multi-agent discussion. Additionally, it evaluates whether the generated content aligns with the topic and proposed outlines. Based on these criteria, the supervisor provides an evaluation and passes the feedback to the next role. Depending on the specific needs, the next role can be the writer, reader, or editor.
@@
Table 2: Comparison of article generation performance for textual content. We evaluate content quality, informativeness, reliability, and engagement under three input modalities (Text-only, Image-only, and Image-Text) on our WikiSeek benchmark.
Writer viewpoint. From the writer's viewpoint, the primary focus is on the multi-agent knowledge exploration and article generation stages. The writer evaluates whether the generated content maintains coherence, ensures engagement, verifies factual accuracy, and upholds logical consistency. Based on these assessments, the writer provides targeted improvement suggestions.
For instance, to enhance coherence, it may be recommended to rearrange sentences or add transitional words and phrases. To improve readability, it might suggest simplifying complex concepts. Finally, the writer responds with a set of targeted and refined suggestions.
@@
Editor viewpoint. After inserting images into the generated article, there may still be discrepancies or inconsistencies between the visual content and the corresponding textual descriptions. To address this, our framework sends the images along with their related text segments to the multi-perspective self-reflection module. This module evaluates the alignment and coherence between the images and their accompanying texts from an editorial viewpoint. It provides targeted suggestions, such as refining the image captions, adjusting image placement, or enhancing textual explanations to better reflect visual content. This final step ensures enhanced relevance, coherence, and readability between visual and textual elements.